OpenStudy (anonymous):

jimmyrep, Help Three cards are drawn from a standard deck of cards. Find each probability.

OpenStudy (anonymous):

1. P(3 hearts), if replacement occurs 2. P(3 hearts), if no replacement occurs

OpenStudy (anonymous):

Assuming replacement means putting the card back in the deck then 1 is 13/52 or 1/4 per card wile in 2 the fraction will reduce top and bottom by 1 after each card.

OpenStudy (anonymous):

estudier,can you check my answer 1) \[\frac{13}{52}*\frac{13}{52}*\frac{13}{52}\]?

OpenStudy (anonymous):

A quarter cubed is 1/64.